# Machine learning Hubbard parameters with equivariant neural networks

### v1

@@ Abstract
Density-functional theory with extended Hubbard functionals (DFT+$U$+$V$) provides a robust framework to accurately describe complex materials containing transition-metal or rare-earth elements. It does so by mitigating self-interaction errors inherent to semi-local functionals which are particularly pronounced in systems with partially-filled $d$ and $f$ electronic states. However, achieving accuracy in this approach hinges upon the accurate determination of the on-site $U$ and inter-site $V$ Hubbard parameters. In practice, these are obtained either by semi-empirical tuning, requiring prior knowledge, or, more correctly, by using predictive but expensive first-principles calculations. Here, we present a machine learning model based on equivariant neural networks which uses atomic occupation matrices as descriptors, directly capturing the electronic structure, local chemical environment, and oxidation states of the system at hand. We target here the prediction of Hubbard parameters computed self-consistently with iterative linear-response calculations, as implemented in density-functional perturbation theory (DFPT), and structural relaxations. Remarkably, when trained on data from 11 materials spanning various crystal structures and compositions, our model achieves mean absolute relative errors of 3% and 5% for Hubbard $U$ and $V$ parameters, respectively. By circumventing computationally expensive DFT or DFPT self-consistent protocols, our model significantly expedites the prediction of Hubbard parameters with negligible computational overhead, while approaching the accuracy of DFPT. Moreover, owing to its robust transferability, the model facilitates accelerated materials discovery and design via high-throughput calculations, with relevance for various technological applications.
@@ INTRODUCTION
A fundamental tool in investigating compounds involving transition-metal (TM) and rare-earth (RE) compounds is density-functional theory (DFT)[1, 2] a cornerstone for first-principles simulations in physics, chemistry, and materials science. In practical applications, DFT necessitates approximations to the exchange-correlation (xc) functional, with the local spin-density approximation (LSDA) and spin-polarized generalized-gradient approximation ($\sigma$-GGA) being the most prevalent choices. However, these approximations yield unsatisfactory outcomes for various properties of TM and RE compounds, primarily due to significant self-interaction errors (SIEs) [3, 4, 5] that are particularly pronounced for localized $d$ and $f$ electrons. To address these challenges, more accurate approaches surpassing the limitations of “standard DFT” have been devised. Noteworthy among these are Hubbard-corrected DFT (so-called DFT+$U$ [6, 7, 8] and its extension DFT+$U$+$V$[9, 10, 11] whose role in addressing SIEs, rather than correlation errors, was first pointed out in Ref. 4), meta-GGA functionals, [12, 13, 14] and hybrid functionals. [15, 16, 17] While these methods offer valuable insights, each comes with inherent limitations and challenges, as discussed in, e.g., Ref. 18.
[…]
Figure 1: (a) Protocol for the self-consistent calculation of Hubbard parameters using DFPT[19] $U_{\mathrm{in}}$ and $V_{\mathrm{in}}$ represent the input Hubbard parameters, while $U_{\mathrm{out}}$ and $V_{\mathrm{out}}$ denote the output parameters, with $\Delta$ representing the convergence threshold. $U_{\mathrm{SC}}$ and $V_{\mathrm{SC}}$ are the final SC Hubbard parameters. (b) Convergence of the Hubbard $U$ parameter for Mn-3d states in LiMnPO4 using the self-consistent protocol.[18] The inset displays the crystal structure of the material, where arrows indicate the spin direction, and Li atoms are depicted in grey, O in red, Mn in violet, and P in yellow.
Hubbard-corrected DFT in particular stands out for its greater accuracy with only a marginal increase in computational cost over standard DFT functionals. [20] In the DFT+$U$+$V$ scheme, a corrective Hubbard energy $E_{U+V}$ is introduced alongside the approximate DFT energy $E_{\mathrm{DFT}}$:
@@
In the simplified rotationally-invariant formulation, [8] the extended Hubbard correction energy for a manifold with angular momentum $\ell$ takes the form: [21]
$E_{U+V}=\frac{1}{2}\sum_{I\sigma}U^{I}\mathrm{Tr}\left[\mathbf{n}_{\ell}^{I}\left(\mathbf{1}-\mathbf{n}_{\ell}^{I\sigma}\right)\right]\\ -\frac{1}{2}\sum_{I}\sum_{J(J\neq I)}^{*}\sum_{\sigma}V^{IJ}\mathrm{Tr}\left[\mathbf{n}^{IJ\sigma}_{\ell}\mathbf{n}^{JI\sigma}_{\ell}\right]\,,$ (2)
where $I$ and $J$ are atomic site indices, while $m$ and $m^{\prime}$ denote magnetic quantum numbers associated with a specific angular momentum. $U^{I}$ and $V^{IJ}$ are effective on-site and inter-site Hubbard parameters, respectively. The asterisk in the sum signifies that, for each atom $I$, the index $J$ encompasses all its neighbors up to a given distance. The generalized occupation matrices $\mathbf{n}^{IJ\sigma}_{\ell}$ are derived from the projection of the Kohn-Sham (KS) states onto localized atom-centered orbitals $\phi^{I}_{m}(\mathbf{r})$ (Hubbard projector functions) of neighboring atoms:
@@
where $v$ and $\sigma$ represent the band and spin labels of the KS wavefunctions $\psi^{\sigma}_{v,\mathbf{k}}(\mathbf{r})$, respectively, $\mathbf{k}$ denotes points in the first Brillouin zone (BZ), and $f^{\sigma}_{v,\mathbf{k}}$ are the occupations of the KS states. The two terms in eq. 2 — proportional to on-site $U^{I}$ and inter-site $V^{IJ}$ — counteract one another. The on-site term promotes localization on atomic sites, suppressing hybridization with neighbors, while the inter-site term favors hybridized states with components on neighboring atoms. Consequently, the values of $U^{I}$ and $V^{IJ}$ are critical to optimizing the extent of the localization and hybridization within Hubbard-corrected DFT. However, these parameters are not known a priori and must be determined in some way. In passing we note that for the sake of simplicity, hereafter we drop the superscripts $I$ and $J$ in the notations of Hubbard parameters and occupation matrices, unless required for clarity.
@@
Hubbard $U$ can be fit semi-empirically to reproduce a target property from experimental data [22, 23, 24, 25, 26] or from other advanced first-principles methods (e.g. $GW$ [27] or hybrids [15, 16, 17]); however, this approach has many limitations. Fitted parameters are not guaranteed to, and often do not, accurately predict properties other than that used in fitting, and calibrating a single Hubbard $U$ to multiple properties is non-trivial, requiring advanced algorithms like Bayesian optimization (BO)[28, 29] This method also precludes materials discovery, where properties are by definition unknown, and targeting results of advanced first-principles methods, like DFT with hybrid functionals, inherits the limitations of those methods. [30, 31] Moreover, the inter-site $V$ parameters, which are necessary to properly describe materials with pronounced covalent interactions, are difficult to fit semi-empirically due to the high-dimensional regression procedure they would require. An attractive alternative is computing Hubbard parameters using first-principles methods such as constrained DFT (cDFT)[32, 33, 34, 35, 36, 37, 38, 39, 40] Hartree-Fock-based approaches, [41, 42, 43, 44, 10, 11] and the constrained random-phase approximation (cRPA)[45, 46, 47, 48] These methods do not rely on data from experiments or advanced simulation methods and, moreover, are able to provide values not only for $U$ but also for $V$. However, first-principles approaches are considerably more computationally expensive than DFT +$U$(+$V$) ground-state calculations themselves, making their application feasible but demanding for large systems or high-throughput studies.
@@
The linear-response (LR) formulation of cDFT (LR-cDFT[21] has witnessed popularity due to its simplicity and accuracy; however, it demands computationally expensive supercell calculations. A recent reformulation of LR-cDFT in terms of DFPT [49, 19] significantly reduces the computational burden for determining Hubbard parameters by replacing cumbersome supercell calculations with faster unit-cell calculations and by leveraging symmetries that further diminishes the computational cost by reducing the number of perturbations in reciprocal space (see Sec. S1 in the supplemental information (SI)). Its physical rationale relies on heuristically imposing piecewise linearity of the energy of the system as a function of the population of the Hubbard manifold. [21] Despite the numerous successful applications of DFPT in computing Hubbard parameters, [50, 51, 18, 52, 53, 54, 55, 56, 57] this approach introduces a significant overhead compared to DFT +$U$+$V$ ground-state calculations. Moreover, it has been shown that jointly optimizing Hubbard parameters and the crystal structure, rather than relying on the equilibrium geometry obtained using (semi-)local functionals, can significantly improve the accuracy of the final properties of interest. [58] This is due in part to taking into account the geometry dependence of $U\{\mathbf{R}\}$ and thus the contribution $dU/d\mathbf{R}$ to the Hellmann-Feynman forces. [59] To do so, a self-consistent procedure combining DFPT and structural optimizations (fig. 1(a)) can be used (see Sec. S2 in the SI). [60, 19] Figure 1(b) shows a typical self-consistent determination of Hubbard parameters for LiMnPO4 following this procedure, where each iteration costs approximately one order of magnitude more computational time than a Hubbard-corrected DFT electronic ground-state calculation. Consequently, the adoption of acceleration techniques is highly sought after to expedite the determination of Hubbard parameters while providing a level of accuracy very close to that of DFPT and, prospectively, a simple estimate of $dU/d\mathbf{R}$ and $dV/d\mathbf{R}$.
@@
The application of machine learning (ML) algorithms to predict Hubbard parameters is gaining popularity, as exemplified by a recent work, Ref. 61, which employs a random forest regression model. Ref. 61 uses crystal structure parameters, such as bond lengths and angles (among others), as descriptors for an ML model trained on a database of Mn oxides. In this study, the training data consist of $U$ values fitted using Bayesian optimization to replicate band gaps and band structures obtained from HSE calculations — a concept initially proposed in the earlier work of Ref. 28 within the context of DFT+$U$ and subsequently extended to the DFT+$U$+$V$ framework. [62] The resulting ML model predicts $U$ values such that subsequent DFT+$U$ calculations using these values yield band gaps that are $2-3$ times smaller than the reference HSE values, while the band structures appear qualitatively similar. Despite the significance of these efforts in applying ML to the prediction of Hubbard parameters, numerous critical issues persist, such as the accuracy of the reference training data (HSE often provides unreliable band gaps for solids [30, 31]), the choice of the target properties for the training data (fitting $U$ to reproduce band gaps is questionable), the importance of self-consistency between Hubbard parameters and the crystal structure, as well as the inclusion of inter-site $V$ which has been neglected thus far, to name a few. Addressing these points is highly relevant and necessitates further investigation, which serves as the motivation for this work.
@@
Here, we introduce a novel ML approach based on equivariant neural-networks (ENNs) which aims to replace computationally demanding first-principles DFPT calculations of Hubbard $U$ and $V$ parameters while providing a negligible loss of accuracy for the vast majority of practical applications. Crucially, the model employs: 1) atomic occupation matrices within the DFT+$U$+$V$ framework as descriptors of the geometry, electronic structure (e.g. oxidation states (OSs)) and local chemical environments in materials, as well as 2) DFPT-based Hubbard parameters, and 3) interatomic distances. This model is general and can be applied to materials with ionic, covalent, and mixed ionic-covalent interactions. It is trained on all intermediate Hubbard parameters and occupation matrices obtained during the self-consistent cycle and directly provides the final self-consistent values of $U$ and $V$, thus also bypassing intermediate structural optimizations. The utilization of ENNs facilitates the exploitation of the inherent O(3) group structure of the occupation matrices, ensuring excellent model performance even with scarce training data. Equivariant models have demonstrated state-of-the-art accuracy and transferability in ML interaction potentials, [63, 64] while our work is the first to incorporate electronic-structure degrees of freedom as explicit features in solids.
@@ RESULTS
The goal of our ML approach is to replace the self-consistent procedure represented in fig. 1(a) and provide the final Hubbard parameters using as input results from an initial DFT(+$U$+$V$) calculation. Conventional ML methods for atomistic systems primarily use the ionic structures as inputs (see e.g. Refs. 65, 66, 67, 68), side-stepping the need to explicitly calculate or even consider the electronic structure. However, as a consequence such models may not be particularly sensitive to changes in electronic structure that only lead to subtle changes in local atomic geometry, as can happen during a change of OS. To overcome this limitation, our models take as input the on-site occupation matrices $\mathbf{n}^{\sigma}_{\ell}\equiv\mathbf{n}_{\ell}^{II\sigma}$ [see eq. 3, $I=J$], which describe the local electronic structure around an atom, including the OS which is reflected in the occupation matrix eigenvalues. [69] This is a particularly compelling choice given the significant variations in Hubbard parameters seen for different OSs of TM elements. [18, 52]
As with any learning task, the ML model should respect the way in which the target physical properties transform under global rotations, translations, reflections and permutations of labels. In the present case, the output $U$ and $V$ values are left unchanged by these transformations, however the entries of the occupation matrix do change under rotation, making $n^{\sigma}_{mm^{\prime}}$ unsuitable for use as inputs to non-symmetry aware models. For this reason, we make use of an equivariant learning model, [70, 71, 72, 73] which has the property that the learned function $f:X\to Y$ obeys the following relation:
@@
where the equivariance is with respect to a group $G$ (in our case SE(3)), and $D_{Y}[g]$ and $D_{X}[g]$ are representations that act on the vector spaces $Y$ and $X$, respectively. For example, applying a rotation to the inputs and then applying $f$ must produce the same outputs as applying the rotation to the outputs of $f$ given the original (unrotated) input. While our model outputs scalars (which are symmetry invariant), by using an equivariant model we ensure that the inputs and all intermediate (hidden) features transform together under group actions. This property has been shown to give state-of-the-art accuracy and transferability, particularly within the domain of ML interaction potentials. [63, 74, 68]
@@
By virtue of being atom-centred, the on-site occupation matrix $\mathbf{n}^{\sigma}_{\ell}$ is naturally invariant to global translations. However, to build a rotationally invariant model, it is convenient to re-express the occupation matrix in terms of the irreducible representations (irreps) of the O(3) group. If we let $D^{(\ell,P)}$ be the irrep with degree $\ell$ and parity $P$, we can express the occupation matrices as the tensor products $D^{(\ell,P)}\otimes D^{(\ell,P)}$, where e.g. $\ell=1$ for $p$ orbitals, and $\ell=2$ for $d$ orbitals. These can then be decomposed into a direct sum of irreps, e.g. for $\ell=1$ and $P=-1$ the occupation matrix is a $3\times 3$, rank-2, tensor $D^{(1,-1)}\otimes D^{(1,-1)}$, that is symmetric, as $\mathbf{n}^{\sigma}_{\ell}=\left(\mathbf{n}^{\sigma}_{\ell}\right)^{\intercal}$. In practice, this decomposition is achieved by applying a change of basis for $\mathbf{n}^{\sigma}_{\ell}$ that transforms it into the irrep basis (see SI S3 for details). At this stage, the remaining symmetry to be addressed is that of label permutation. In the case of the atom labels for the inter-site term, the model should be invariant to permutation of $I$ and $J$, however in our training data the Hubbard $V$ correction is always applied to a $d$-block/$p$-block atom pairs which have occupation matrices of different dimension ($5\times 5$ for $d$ and $3\times 3$ for $p$), thereby making them distinguishable.
@@
$U^{I}$, $V^{IJ}$, $r_{IJ}$ $D^{(0,1)}$
Atomic species $\bigoplus^{N_{s}}D^{(0,1)}$
$\mathbf{n}^{\sigma}_{\mathrm{p}}$ $D^{(0,1)}\oplus D^{(2,1)}$
$\mathbf{n}^{\sigma}_{\mathrm{d}}$ $D^{(0,1)}\oplus D^{(2,1)}\oplus D^{(4,1)}$
Table 1: Irreducible representations of various attributes (inputs and outputs) of the ML model. $U^{I}$ and $V^{IJ}$ are the on-site and inter-site Hubbard parameters, respectively, $r_{IJ}$ is the interatomic distance between species $I$ and $J$, $N_{s}$ is the number of atomic species (one-hot vectors), $\mathbf{n}^{\sigma}_{\mathrm{p}}$ and $\mathbf{n}^{\sigma}_{\mathrm{d}}$ are the occupation matrices for the p and d orbitals, respectively, in the special representation for the ML model [see eqs. 5 and 6]. The irreducible representations of the occupation matrices can be readily obtained using the e3nn [73, 75] library.
[…]
Figure 2: Schematic illustration of the equivariant neural-network ML model for predicting on-site Hubbard $U$ (a) and inter-site Hubbard $V$ (b) parameters. $U_{\mathrm{in}}$ and $V_{\mathrm{in}}$ are the input Hubbard parameters, while $U_{\mathrm{out}}$ and $V_{\mathrm{out}}$ are the outputs. The atomic species enter as one-hot tensors, and $r_{IJ}$ is the interatomic distance between sites $I$ and $J$. $\mathbf{x}^{1}_{\mathrm{d}}$ and $\mathbf{x}^{2}_{\mathrm{d}}$ are the occupation matrices for the $d$ orbitals, while $\mathbf{x}^{1}_{\mathrm{p}}$ and $\mathbf{x}^{2}_{\mathrm{p}}$ for the $p$ orbitals, all in the special representation for the ML model [see eqs. 5 and 6]. Tensors are represented as squares: open for inputs and outputs, and filled for intermediate features.

### v2

@@ Abstract
Density-functional theory with extended Hubbard functionals (DFT+$U$+$V$) provides a robust framework to accurately describe complex materials containing transition-metal or rare-earth elements. It does so by mitigating self-interaction errors inherent to semi-local functionals which are particularly pronounced in systems with partially-filled d and f electronic states. However, achieving accuracy in this approach hinges upon the accurate determination of the on-site $U$ and inter-site $V$ Hubbard parameters. In practice, these are obtained either by semi-empirical tuning, requiring prior knowledge, or, more correctly, by using predictive but expensive first-principles calculations. Here, we present a machine learning model based on equivariant neural networks which uses atomic occupation matrices as descriptors, directly capturing the electronic structure, local chemical environment, and oxidation states of the system at hand. We target here the prediction of Hubbard parameters computed self-consistently with iterative linear-response calculations, as implemented in density-functional perturbation theory (DFPT), and structural relaxations. Remarkably, when trained on data from 12 materials spanning various crystal structures and compositions, our model achieves mean absolute relative errors of 3% and 5% for Hubbard $U$ and $V$ parameters, respectively. By circumventing computationally expensive DFT or DFPT self-consistent protocols, our model significantly expedites the prediction of Hubbard parameters with negligible computational overhead, while approaching the accuracy of DFPT. Moreover, owing to its robust transferability, the model facilitates accelerated materials discovery and design via high-throughput calculations, with relevance for various technological applications.
@@ INTRODUCTION
A fundamental tool in investigating compounds involving transition-metal (TM) and rare-earth (RE) compounds is density-functional theory (DFT)1, 2 a cornerstone for first-principles simulations in physics, chemistry, and materials science. In practical applications, DFT necessitates approximations to the exchange-correlation (xc) functional, with the local spin-density approximation (LSDA) and spin-polarized generalized-gradient approximation ($\sigma$-GGA) being the most prevalent choices. However, these approximations yield unsatisfactory outcomes for various properties of TM and RE compounds, primarily due to significant self-interaction errors (SIEs) 3, 4, 5 that are particularly pronounced for localized d and f electrons. To address these challenges, more accurate approaches surpassing the limitations of “standard DFT” have been devised. Noteworthy among these are Hubbard-corrected DFT (so-called DFT+$U$ 6, 7, 8 and its extension DFT+$U$+$V$9, 10, 11 whose role in addressing SIEs, rather than correlation errors, was first pointed out in Ref. 4), meta-GGA functionals, 12, 13, 14 and hybrid functionals. 15, 16, 17 While these methods offer valuable insights, each comes with inherent limitations and challenges, as discussed in, e.g., Ref. 18. Hubbard-corrected DFT in particular stands out for its greater accuracy with only a marginal increase in computational cost over standard DFT functionals. 19 In the DFT+$U$+$V$ scheme, a corrective Hubbard energy $E_{U+V}$ is introduced alongside the approximate DFT energy $E_{\mathrm{DFT}}$:
@@
In the simplified rotationally-invariant formulation, 8 the extended Hubbard correction energy for a manifold with angular momentum $\ell$ takes the form: 20
$E_{U+V}=\frac{1}{2}\sum_{I\sigma}U^{I}\mathrm{Tr}\left[\mathbf{n}_{\ell}^{II\sigma}\left(\mathbf{1}-\mathbf{n}_{\ell}^{II\sigma}\right)\right]\\ -\frac{1}{2}\sum_{I}\sum_{J(J\neq I)}^{*}\sum_{\sigma}V^{IJ}\mathrm{Tr}\left[\mathbf{n}^{IJ\sigma}_{\ell}\mathbf{n}^{JI\sigma}_{\ell}\right]\,,$ (2)
where $I$ and $J$ are atomic site indices, and $\sigma$ is the spin index. $U^{I}$ and $V^{IJ}$ are effective on-site and inter-site Hubbard parameters, respectively. The asterisk in the sum signifies that, for each atom $I$, the index $J$ encompasses all its neighbors up to a given distance. The generalized occupation matrices $\mathbf{n}^{IJ\sigma}_{\ell}$ are derived from the projection of the Kohn-Sham (KS) states onto localized atom-centered orbitals $\phi^{I}_{m}(\mathbf{r})$ (Hubbard projector functions) of neighboring atoms:
@@
where $v$ represents the band index of the KS wavefunctions $\psi^{\sigma}_{v,\mathbf{k}}(\mathbf{r})$, $m$ and $m^{\prime}$ denote magnetic quantum numbers associated with a specific angular momentum $\ell$, $\mathbf{k}$ denotes points in the first Brillouin zone (BZ), and $f^{\sigma}_{v,\mathbf{k}}$ are the occupations of the KS states. The two terms in eq. 2 — proportional to on-site $U^{I}$ and inter-site $V^{IJ}$ — counteract one another. The on-site term promotes localization on atomic sites, suppressing hybridization with neighbors, while the inter-site term favors hybridized states with components on neighboring atoms. Consequently, the values of $U^{I}$ and $V^{IJ}$ are critical to optimizing the extent of the localization and hybridization within Hubbard-corrected DFT. However, these parameters are not known a priori and must be determined in some way. In passing we note that for the sake of simplicity, hereafter we drop the superscripts $I$ and $J$ in the notations of Hubbard parameters and occupation matrices, unless required for clarity.
@@
Hubbard $U$ can be fit semi-empirically to reproduce a target property from experimental data 21, 22, 23, 24, 25 or from other advanced first-principles methods (e.g. $GW$ 26 or hybrids 15, 16, 17); however, this approach has many limitations. Fitted parameters are not guaranteed to, and often do not, accurately predict properties other than those used for fitting, and calibrating a single Hubbard $U$ to multiple properties is non-trivial, requiring advanced algorithms like Bayesian optimization (BO)27, 28 This method also precludes materials discovery, where properties are by definition unknown, and targeting results of advanced first-principles methods, like DFT with hybrid functionals, inherits the limitations of those methods. 29, 30 Moreover, the inter-site $V$ parameters, which are necessary to properly describe materials with pronounced covalent interactions, are difficult to fit semi-empirically due to the high-dimensional regression procedure they would require. An attractive alternative is computing Hubbard parameters using first-principles methods such as constrained DFT (cDFT)31, 32, 33, 34, 35, 36, 37, 38, 39 Hartree-Fock-based approaches, 40, 41, 42, 43, 10, 11 and the constrained random-phase approximation (cRPA)44, 45, 46, 47 These methods do not rely on data from experiments or advanced simulation methods and, moreover, are able to provide values not only for $U$ but also for $V$. However, first-principles approaches are considerably more computationally expensive than DFT +$U$(+$V$) ground-state calculations themselves, making their application feasible but demanding for large systems or high-throughput studies.
Figure 1: Calculating Hubbard corrections self-consistently using density-functional perturbation theory. a) Protocol for the self-consistent calculation of Hubbard parameters using DFPT48 $U_{\mathrm{in}}$ and $V_{\mathrm{in}}$ represent the input Hubbard parameters, while $U_{\mathrm{out}}$ and $V_{\mathrm{out}}$ denote the output parameters, with $\Delta$ representing the convergence threshold. $U_{\mathrm{SC}}$ and $V_{\mathrm{SC}}$ are the final SC Hubbard parameters. b) Convergence of the Hubbard $U$ parameter for Mn-3d states in LiMnPO4 using the self-consistent protocol.18 The inset displays the crystal structure of the material, where arrows indicate the spin direction, and Li atoms are depicted in grey, O in red, Mn in violet, and P in yellow.
@@
The linear-response (LR) formulation of cDFT (LR-cDFT20 has witnessed popularity due to its simplicity and accuracy; however, it demands computationally expensive supercell calculations. A recent reformulation of LR-cDFT in terms of DFPT 49, 48 significantly reduces the computational burden for determining Hubbard parameters by replacing cumbersome supercell calculations with faster unit-cell calculations and by leveraging symmetries that further diminishes the computational cost by reducing the number of perturbations in reciprocal space (see Sec. S1 in the supplemental information (SI)). Its physical rationale relies on heuristically imposing piecewise linearity of the total energy of the system as a function of the population of the Hubbard manifold. 20 Despite the numerous successful applications of DFPT in computing Hubbard parameters, 50, 51, 18, 52, 53, 54, 55, 56, 57 this approach introduces a significant overhead compared to DFT +$U$+$V$ ground-state calculations. Moreover, it has been shown that jointly optimizing Hubbard parameters and the crystal structure, rather than relying on the equilibrium geometry obtained using (semi-)local functionals, can significantly improve the accuracy of the final properties of interest. 58 This is due in part to taking into account the geometry dependence of $U\{\mathbf{R}\}$ and thus the contribution $dU/d\mathbf{R}$ to the Hellmann-Feynman forces. 59 To do so, a self-consistent procedure combining DFPT and structural optimizations (fig. 1) can be used (see Sec. S2 in the SI). 60, 48 Figure 1 shows a typical self-consistent determination of Hubbard parameters for LiMnPO4 following this procedure, where each iteration costs approximately one order of magnitude more computational time than a Hubbard-corrected DFT electronic ground-state calculation. Consequently, the adoption of acceleration techniques is highly sought after to expedite the determination of Hubbard parameters while providing a level of accuracy very close to that of DFPT and, prospectively, a simple estimate of $dU/d\mathbf{R}$ and $dV/d\mathbf{R}$.
@@
Using machine learning (ML) to accelerate parts of a DFT workflow is becoming increasingly common, 61, 62 including applications to predicting Hubbard parameters. A recent example is the work of Cai et al. 63, in which they use crystal structure parameters, such as bond lengths and angles (among others), as descriptors for a random forest regression model trained on a database of Mn oxides. Their training data consist of $U$ values fitted using Bayesian optimization to replicate band gaps and band structures obtained from HSE hybrid functional calculations — a concept initially proposed in the earlier work of Ref. 27 within the context of DFT+$U$ and subsequently extended to the DFT+$U$+$V$ framework. 64 The resulting ML model predicts $U$ values such that subsequent DFT+$U$ calculations using these values yield band gaps that are $2-3$ times smaller than the reference HSE values, while the band structures appear qualitatively similar. Despite the significance of these efforts in applying ML to the prediction of Hubbard parameters, numerous critical issues persist, such as the accuracy of the reference training data (HSE often provides unreliable band gaps for solids 29, 30), the choice of the target properties for the training data (fitting $U$ to reproduce band gaps is questionable), the importance of self-consistency between Hubbard parameters and the crystal structure, as well as the lack of inter-site $V$, to name a few. Addressing these points is highly relevant and necessitates further investigation, which serves as the motivation for this work.
@@
Here, we introduce a novel ML approach based on equivariant neural-networks (ENNs) which aims to replace computationally demanding first-principles DFPT calculations of Hubbard $U$ and $V$ parameters while providing a negligible loss of accuracy for the vast majority of practical applications. Crucially, the model employs: 1) atomic occupation matrices within the DFT+$U$+$V$ framework as descriptors of the geometry, electronic structure (e.g. oxidation states (OSs)) and local chemical environments in materials, as well as 2) DFPT-based Hubbard parameters, and 3) interatomic distances. This model is general and can be applied to materials with ionic, covalent, and mixed ionic-covalent interactions. It is trained on all intermediate Hubbard parameters and occupation matrices obtained during the self-consistent cycle and directly provides the final self-consistent values of $U$ and $V$, thus also bypassing intermediate structural optimizations. The utilization of ENNs facilitates the exploitation of the inherent O(3) group structure of the occupation matrices, ensuring excellent model performance even with scarce training data. Equivariant models have demonstrated state-of-the-art accuracy and transferability in ML interaction potentials, 65, 66 while our work is the first to incorporate electronic-structure degrees of freedom as explicit features in solids.
@@ RESULTS
The goal of our ML approach is to replace the self-consistent procedure represented in fig. 1 and provide the final Hubbard parameters using as input results from an initial DFT(+$U$+$V$) calculation. Conventional ML methods for atomistic systems primarily use the ionic structures as inputs (see e.g. Refs. 67, 68, 69, 70), side-stepping the need to explicitly calculate or even consider the electronic structure. However, as a consequence such models may not be particularly sensitive to changes in electronic structure that only lead to subtle changes in local atomic geometry, as can happen during a change of OS. To overcome this limitation, our models take as input the on-site occupation matrices $\mathbf{n}^{\sigma}_{\ell}\equiv\mathbf{n}_{\ell}^{II\sigma}$ [see eq. 3, $I=J$], which describe the local electronic structure around an atom, including the OS which is reflected in the occupation matrix eigenvalues. 71 This is a particularly compelling choice given the significant variations in Hubbard parameters seen for different OSs of TM elements. 18, 52 We emphasize that relying solely on the trace of the occupation matrix as a descriptor may not adequately distinguish between different OSs and chemical environments.
As with any learning task, the ML model should respect the way in which the target physical properties transform under global rotations, translations, reflections and permutations of labels. In the present case, the output $U$ and $V$ values are left unchanged by these transformations, however the entries of the occupation matrix do change under rotation, making $n^{\sigma}_{mm^{\prime}}$ unsuitable for use as inputs to non-symmetry aware models. For this reason, we make use of an equivariant learning model, 72, 73, 74, 75 which has the property that the learned function $f:X\to Y$ obeys the following relation:
@@
where the equivariance is with respect to a group $G$ (in our case SE(3)), and $D_{Y}[g]$ and $D_{X}[g]$ are representations that act on the vector spaces $Y$ and $X$, respectively. For example, applying a rotation to the inputs and then applying $f$ must produce the same outputs as applying the rotation to the outputs of $f$ given the original (unrotated) input. While our model outputs scalars (which are symmetry invariant), by using an equivariant model we ensure that the inputs and all intermediate (hidden) features transform together under group actions. This property has been shown to give state-of-the-art accuracy and transferability, particularly within the domain of ML interaction potentials. 65, 76, 70
@@
By virtue of being atom-centred, the on-site occupation matrix $\mathbf{n}^{\sigma}_{\ell}$ is naturally invariant to global translations. However, to build a rotationally invariant model, it is convenient to re-express the occupation matrix in terms of the irreducible representations (irreps) of the O(3) group. If we let $\Gamma^{(\ell,P)}$ be the irrep with degree $\ell$ and parity $P$, we can express the occupation matrices as the tensor products $\Gamma^{(\ell,P)}\otimes\Gamma^{(\ell,P)}$, where e.g. $\ell=1$ for p orbitals, and $\ell=2$ for d orbitals. These can then be decomposed into a direct sum of irreps, e.g. for $\ell=1$ and $P=-1$ the occupation matrix is a $3\times 3$, rank-2, tensor $\Gamma^{(1,-1)}\otimes\Gamma^{(1,-1)}$, that is symmetric, as $\mathbf{n}^{\sigma}_{\ell}=\left(\mathbf{n}^{\sigma}_{\ell}\right)^{\intercal}$. In practice, this decomposition is achieved by applying a change of basis for $\mathbf{n}^{\sigma}_{\ell}$ that transforms it into the irrep basis (see Supplementary Section 3 for details). At this stage, the remaining symmetry to be addressed is that of label permutation. In the case of the atom labels for the inter-site term, the model should be invariant to permutation of $I$ and $J$, however in our training data the Hubbard $V$ correction is always applied to a d-block/p-block atom pairs which have occupation matrices of different dimension ($5\times 5$ for d and $3\times 3$ for p), thereby making them distinguishable.
@@
$U^{I}$, $V^{IJ}$, $r_{IJ}$ $\Gamma^{(0,1)}$
Atomic species $\bigoplus^{N_{s}}\Gamma^{(0,1)}$
$\mathbf{n}^{\sigma}_{\mathrm{p}}$ $\Gamma^{(0,1)}\oplus\Gamma^{(2,1)}$
$\mathbf{n}^{\sigma}_{\mathrm{d}}$ $\Gamma^{(0,1)}\oplus\Gamma^{(2,1)}\oplus\Gamma^{(4,1)}$
Table 1: Irreducible representations of various attributes (inputs and outputs) of the ML model. $U^{I}$ and $V^{IJ}$ are the on-site and inter-site Hubbard parameters, respectively, $r_{IJ}$ is the interatomic distance between species $I$ and $J$, $N_{s}$ is the number of atomic species (one-hot vectors), $\mathbf{n}^{\sigma}_{\mathrm{p}}$ and $\mathbf{n}^{\sigma}_{\mathrm{d}}$ are the occupation matrices for the p and d orbitals, respectively, in the special representation for the ML model [see eqs. 5 and 6]. The irreducible representations of the occupation matrices can be readily obtained using the e3nn 75, 77 library.